The Fermi enrgy of copper is 7 eV. (use h = 6.6 ×10−34JHz−1 , 1eV = 1.602×10−19J, me=9.1×10−31 kg )
The Fermi momentum of electron in copper is ________

16.48×10−25 kg m/s


13.38×10−25 kg m/s


12.48×10−25 kg m/s


14.28×10−25 kg m/s
1
Expert's answer
2021-09-23T11:26:41-0400

The Fermi momentum is given as follows (see https://en.wikipedia.org/wiki/Fermi_energy):



"p_F = \sqrt{2E_Fm_e}"

where "E_F = 7eV = 7\cdot 1.602×10^{−19}J \approx 1.121\times 10^{-18}J" is the Fermi energy, and "m_e = 9.1\u00d710^{\u221231} kg" is the mass of electron. Thus, obtain:


"p_F = \sqrt{2\cdot 1.121\times 10^{-18}\cdot 9.1×10^{−31}} \approx 14.28\times 10^{-25}kg\cdot m/s"



Answer. 14.28×10−25 kg m/s.